OpenStudy (anonymous):

Write the product as a trinomial. (k – 3)(k – 1) A. k2– 3 B. k2 – 4k + 3 C. k2 – 4k – 3 D. k2 + 4k + 3

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

Can you expand (k – 3)(k – 1) ? I will give you an example (n – 2)(n – 6) = n(n – 6) -2(n-6) = n^2-6n -2n +12=n^2-8n-12 is the answer to the example. No to start you off, tell me k(k – 3) = ? (-1)(k-3) = ?

OpenStudy (nurali):

(k – 3)(k – 1) k^2-k-3k+3 k^2-4k+3